OpenStudy (jesuslover11):

What is the area of a rectangle with a perimeter of 40 m and a length of 8 m? a) 30 m b) 96 m c) 86 m

OpenStudy (anonymous):

B

OpenStudy (anonymous):

A = Base*Width

OpenStudy (jesuslover11):

Thank u!! wait a or b?

OpenStudy (anonymous):



OpenStudy (anonymous):

8 + 8 + x + x = 40 16 + 2x = 40 24 = 2x 12 = x A = 12 *8 A= 96